OpenStudy (themathasker):

Sally and Brianna are baking cakes, Sally need 2 1/3 cups of flours for her recipe, and Brianna needs 1 3/4 for hers. They have 4 cups of flour, do they have enough for both recipes?

OpenStudy (anonymous):

So you need to make a comparison. First, you need to add the fractions. To add the fractions they must have the same denominator. So 2 1/3 becomes 2 4/12 and 1 3/4 becomes 1 9/12. Now make them mixed fractions easier to add. So you have 28/12 + 21/12= 49/12 now convert by dividing. So 49/12 = 4 1/12. so no they do not

OpenStudy (themathasker):

Thank you c: